OpenStudy (accidentalaichan):

Hm, I seem to be a stick in the mud with this one: What is the slope of the line that passes through (0,1) and is parallel to y = 5x + 3?

jimthompson5910 (jim_thompson5910):

The slope of y = 5x + 3 is m = ______ (fill in the blank)

OpenStudy (accidentalaichan):

m=5?

jimthompson5910 (jim_thompson5910):

so the slope of any line parallel to y = 5x + 3 is also 5 this is because if two lines are parallel, then their slopes are equal

jimthompson5910 (jim_thompson5910):

we know that m = 5 (parallel slope) and the given point on this parallel line is (0,1) so that means that when x = 0, y = 1

jimthompson5910 (jim_thompson5910):

y = mx + b y = 5x + b ... plug in the parallel slope m = 5 1 = 5(0) + b ... plug in the given point (0,1) now solve for b and tell me what you get

OpenStudy (accidentalaichan):

1 = 0 + b 1 = b

jimthompson5910 (jim_thompson5910):

so m = 5, b = 1

jimthompson5910 (jim_thompson5910):

y = mx+b y = 5x + 1

jimthompson5910 (jim_thompson5910):

that's the equation of the parallel line that goes through (0,1)

OpenStudy (accidentalaichan):

Thanks so much!

jimthompson5910 (jim_thompson5910):

you're welcome